quickstudent:

A triangle has a perimeter of 50. If two of the side lengths are equal and the third side is five more than the equal sides, what is the length of the third side? 15 23 20 35

Hero:

Let x represent the length of the 1st side and y represent the length of the 2nd side and z represent the length of the 3rd side Then x + y + z = 50 and x = y since the first couple lengths are equal and z = x + 5 represents the length of the 3rd side since it is five more than both x and y. so x + x + (x + 5) = 50 In short, 3x + 5 = 50 Solve for x. Take the result you find for x and add five more to get the length of the third side.